solve power series solution


y'' - (x+1) y' -y = 0



1
Expert's answer
2021-06-14T11:13:09-0400

"y=\sum c_nx^n \newline
y'=\sum nc_nx^{n-1} \newline
y''=\sum n(n-1)c_nx^{n-2} \newline
\text{substitute the above value in the given equation, we get}\newline

\implies \sum_{n=2}^\infty n(n-1)c_nx^{n-2}-(x+1)\sum_{n=1}^\infty nc_nx^{n-1}-\sum_{n=0}^\infty c_nx^n=0 \newline

\implies
\sum_{n=0}^\infty (n+2)(n+1)c_{n+2}x^{n}-\sum_{n=1}^\infty nc_{n}x^{n}-\sum_{n=0}^\infty (n+1)c_{n+1}x^{n}
-\sum_{n=0}^\infty c_nx^n=0 \newline

\implies
2c_2-c_1-c_0+\sum_{n=1}^\infty (n+2)(n+1)c_{n+2}x^{n}-\sum_{n=1}^\infty nc_{n}x^{n}-\sum_{n=1}^\infty (n+1)c_{n+1}x^{n}
-\sum_{n=1}^\infty c_nx^n=0 \newline

\implies
2c_2-c_1-c_0+\sum_{n=1}^\infty
[(n+2)(n+1)c_{n+2}- nc_n-(n+1)c_{n+1}-c_n]x^n=0 \newline

\text{We get,}\newline
2c_2-c_1-c_0=0\\
\text{and
the recurrence relation is,}\\
(n+2)(n+1)c_{n+2}- nc_n-(n+1)c_{n+1}-c_n=0\\
(n+2)(n+1)c_{n+2}- (n+1)c_{n+1}-(n+1)c_n=0\\
c_{n+2}=\frac{(n+1)c_{n+1}+(n+1)c_n}{
(n+2)(n+1)}\\
c_{n+2}=\frac{c_{n+1}+c_n}{(n+2)}\\

\text{Thus, the solution is given by }\newline
y=c_0y1+c_1y_2.\\
Let c_0=1, c_1=0, \implies y=y_1.\\
Then, c_2=1/2, c_3=3/4.\\
Therefore, y_1=1+\frac{1}{2}x^2+\frac{1}{4}x^3+•••.\\
Similarly,\newline
 c_0=0, c_1=1, \implies y=y_2.\newline
Then, c_2=1/2, c_3=3/4.\newline
Therefore,
y_2=x+\frac{1}{2}x^2+\frac{3}{4}x^3+•••."